# DiT4Edit: Diffusion Transformer for Image Editing

### v1

@@ 1 Introduction
A primary challenge in text-driven image editing is maintaining the consistency between the source and target images. Earlier approaches (Choi et al. 2021) (Kawar et al. 2023a) (Zhang et al. 2023b) often relied on fine-tuning diffusion models to address this issue. However, these methods typically require considerable time and computation resources, which limits their practical applicability. Recent approaches often utilize DDIM (Song, Meng, and Ermon 2020) inversion to obtain latent maps and then control the attention mechanism in diffusion models for real image editing (Mokady et al. 2023) (Cao et al. 2023). However, the consistency of the edited images depends heavily on the invertibility of the DDIM inversion process. Although some efforts have focused on optimizing this inversion (Ju et al. 2024) (Dong et al. 2023) for better results, the editing framework still relies on too many timesteps (e.g., 50 steps).
@@ 2.1 Text-to-Image Generation
Prior to the emergence of diffusion models, the successful application of Transformer architecture (Vaswani et al. 2017) to image recognition with Vision Transformer (ViT) by Dosovitskiy et al. (Dosovitskiy et al. 2020), and for high-resolution image synthesis with Taming Transformers by Esser et al. (Esser, Rombach, and Ommer 2021), demonstrated the immense potential of Transformers in visual generative tasks. Subsequent methods like SPADE (Park et al. 2019) and time-lapse synthesis by Nam et al. (Nam et al. 2019) advanced the field towards more complex image synthesis tasks. These methods provided richer context awareness and temporal dimension control. The introduction of Denoising Diffusion Probabilistic Models (DDPMs) by Ho et al. (Ho, Jain, and Abbeel 2020) and DALL·E by Ramesh et al. (Ramesh et al. 2021) marked significant leaps forward, achieving breakthroughs in image quality, controllability, and diversity. Designs and applications of diffusion-based methods can be classified by tasks such as controllable generation, stylization, and quality improvement. Nichol et al. (Nichol et al. 2021) explored text-conditional image generation using diffusion models, finding classifier-free guidance to produce realistic images closely aligned with human expectations. Stylization techniques, such as InST (Zhang et al. 2023a) and DiffStyler (Huang et al. 2024), and quality enhancement tools like eDiff-I (Balaji et al. 2022) and ScaleCrafter (He et al. 2023) push the limits of image resolution and fidelity. Moreover, these T2I models have also driven the development of a series of video generation applications (Ma et al. 2022a) (Ma et al. 2023) (Ma et al. 2022b) (Ma et al. 2024c) (Chen et al. 2024). Specifically, Chen et al. (Chen et al. 2023) present PIXART-$\alpha$, a Transformer-based diffusion model that significantly reduces training time and costs while maintaining competitive image quality.
@@ 2.2 Interactive Image Editing
Image editing encompasses scenarios like iterative generation, collaborative creation, and image inpainting. Research has focused on decoupling high-level concepts and low-level styles within deep latent structures to improve diffusion-based models’ performance in tasks such as content editing (detail control) (Kawar et al. 2023b), style transfer (Brack et al. 2022), and textual inversion (Gal et al. 2022). Diffusion-based methods have revolutionized content editing, offering enhanced photorealism and greater control in text-guided synthesis and manipulation of images. Innovations like ILVR (Choi et al. 2021) and ControlNet (Zhang, Rao, and Agrawala 2023) enable precise control over image attributes, while methods like GLIDE (Nichol et al. 2021) and InstructPix2Pix (Brooks, Holynski, and Efros 2023) showcase sophistication in content editing and adaptive response to textual prompts. Hertz et al. (Hertz et al. 2022) introduced a framework for image editing through textual prompts, highlighting the role of cross-attention layers in mapping text to image layout, enabling precise local and global edits. Kumari et al. (Kumari et al. 2023) proposed Custom Diffusion, a method for incorporating user-defined concepts into text-to-image diffusion models, allowing rapid adaptation to new concepts. Brooks et al. (Brooks, Holynski, and Efros 2023) presented InstructPix2Pix, a model that interprets human-written instructions to edit images accurately without needing per-example fine-tuning. Finally, Parmar et al. (Parmar et al. 2023) addressed content preservation in image-to-image translation with pix2pix-zero, ensuring input image content remains intact while facilitating superior performance in both real and synthetic image editing without additional training. Although there have been significant advancements in image editing using diffusion models, these existing attempts of image editing are still bounded by the pretrained generative prior of UNet. Compared with UNet-based diffusion model, DiT is more scalable and has more succinct architecture while DiT’s application in image editing is still under-explored.
@@ 3 Methodology
Our proposed framework aims to achieve high-quality image editing for various sizes based on a diffusion transformer. Our method is the first editing strategy based on a pre-trained text-to-image transformer-based diffusion model, e.g., PIXART-$\alpha$  (Chen et al. 2023). With our approach, users can achieve better editing results compared to existing UNet-based methods by providing a target prompt. In this section, we first introduce the latent diffusion models and DPM inversion. Then we illustrate the superiority of transformer-based denoising in image editing tasks. Finally, we discuss the implementation details of our editing method.
@@ 3.1 Preliminaries: Latent Diffusion Models
Latent Diffusion Models. The Latent Diffusion Model (LDM) (Rombach et al. 2022) proposes an image generation method with a denoising process in the latent space $\mathcal{Z}$. Specifically, it uses an encoder $\mathcal{E}$ to encode a pixel image $x$ into low-resolution latents $z=\mathcal{E}(x)$. During training, the model optimizes a denoising UNet $\epsilon_{\theta}$ by removing artificial noise, conditioned on both text prompt embedding $y$ and current image sample $z_{t}$, which is a noisy sample of $z_{0}$ at step $t\in[0,T]$:
@@ 3.3 Diffusion Transformer-based Image Editing
DPM-Solver inversion. As we discussed earlier, using a high-order DPM-Solver (e.g. DPM-Solver++), can effectively improve the sampling speed. To approximate the integral term $\int_{\lambda_{s}}^{\lambda_{t}}e^{-\lambda}\boldsymbol{x}_{\theta}\mathrm{d}\lambda$  in equation Eq. 4, given the $x_{t_{i-1}}$ at time $t_{t_{i-1}}$, using the Taylor expansion at $\lambda_{t_{i-1}}$, and the DPM-Solver++ can obtain a exact solution value at time $t_{i}$:
@@
The unified control of attention mechanism. In the previous work Prompt to Prompt (P2P) (Hertz et al. 2022), researchers demonstrate that the cross attention layers contain rich semantic information from prompt texts. This finding can edit images through replace the cross attention maps between the source image and target image during the diffusion process. Specifically, the two commonly used text guided cross attention control strategies are cross attention replace and cross attention refine. These two methods ensure the seamless flow of information from the target prompt to the source prompt, thereby guiding the latent map towards the desired direction.
Different from the cross attention, the self-attention mechanism in diffusion transformer is utilized to guide the formation of image layout, a feature that cannot be accomplished by the cross-attention mechanism. As shown in Figure 3, the object and layout information from the prompt are not fully captured in the query vectors of the transformer’s shallow layers but are well-represented in the deeper layers. Moreover, with an increasing number of transformer layers, the query vectors’ ability to capture object details becomes clearer and more specific. This suggests that the transformer’s global attention mechanism is more effective at capturing long-range object information, making DiT particularly advantageous for large-scale deformation and editing of extensive images. This observation suggests that non-rigid editing of images can be achieved by controlling self attention mechanism. In MasaCtrl (Cao et al. 2023), researchers introduced the mutual self attention control mechanism. To be specific, in the early steps of diffusion, the feature in the editing steps $Q_{tar}$, $K_{tar}$, and $V_{tar}$ will be used in self attention calculation to generate an image layout closer to the target prompt, while in the later stages, the feature in the reconstruction steps ––$K_{src}$ and $V_{src}$ will be used to guide the generation of the target image layout closer to the original image.
@@
Table 1: Quantitative comparison results. We compare our model with six prior works: Pix2Pix-Zero (Parmar et al. 2023), PnPInversion (Ju et al. 2024), SDEdit (Meng et al. 2021), IP2P (Brooks, Holynski, and Efros 2023), MasaCtrl (Cao et al. 2023), and InfEdit (Xu et al. 2023), all implemented using official open source code.
@@ 4.1 Implementation Details
For editing tasks involving images with a scale of $512\times 512$ and larger sizes up to $1024\times 2048$, we use pre-trained models PixArt-$\alpha-\text{XL}-512\times 512$ version for the smaller scale and PixArt-$\alpha-\text{XL}-1024\times 1024-\text{MS}$ version for the larger scale (Chen et al. 2023). We conduct editing on both real and generated images. For the real image input, we use DPM-Solver inversion to get the latent noise map. We configured the DPM-Solver with 30 steps, the classifier-free guidance of 4.5, and a patches merging ratio of 0.8. All experiments were conducted on an NVIDIA Tesla A100 GPU.
@@ 4.2 Qualitative Comparison
We evaluate the qualitative performance differences between our proposed DiT4Edit editing framework and six prior baselines, including Pix2Pix-Zero (Parmar et al. 2023), PnPInversion (Ju et al. 2024), SDEdit (Meng et al. 2021), IP2P (Brooks, Holynski, and Efros 2023), MasaCtrl (Cao et al. 2023), and InfEdit (Xu et al. 2023), all implemented using official open source code.
As shown in Figure 5, we compare our method on $512\times 512$ and $1024\times 1024$ images. The first row of the Figure 5 demonstrates our framework has the ability to generate edited images that remain consistent with the original content when editing real $512\times 512$ images, whereas existing methods often alter the background or target details of the original image. Furthermore, the second and third rows of Figure 5 illustrate our experiments with large-scale images and arbitrarily sized images—tasks that previous UNet-based methods struggled to address. The results indicate that our proposed framework effectively handles style and object shape modifications in larger images. In contrast, some state-of-the-art UNet-based methods, despite being capable of performing editing tasks, frequently result in significant alterations and damage to the background and object locations in the source image. Additionally, due to the limitations of the UNet structure, these methods typically generate target images only at a size of $512\times 512$. These findings emphasize the substantial potential of transformer-based diffusion models in large-scale image editing. We also perform the user study for comprehensive comparison. The details of user study can be found in supplementary material.
@@ 4.4 Ablation Study
We conduct a series of ablation studies to demonstrate the effectiveness of the DPM-Solver inversion and patches merging. The results of our ablation experiments on patches merging are presented in Figure 4 and Table 2. Implementing patches merging led to a notable reduction in the editing time for large-sized images while maintaining editing quality comparable to that achieved without patches merging. This indicates that patches merging can significantly enhance the overall performance of image editing frameworks. Furthermore, the ablation experiment results for DPM-Solver and DDIM are illustrated in Figure 7. When comparing the two methods with the same number of inference steps ($T=30$), DPM Solver consistently outperformed DDIM in terms of image editing quality. This demonstrates that our use of the DPM-Solver inversion strategy allows for the generation of superior latent maps, resulting in better editing outcomes within fewer steps.

### v2

@@ 1 Introduction
A primary challenge in text-driven image editing is maintaining the consistency between the source and target images. Earlier approaches (Choi et al. 2021) (Kawar et al. 2023a) (Zhang et al. 2023) often relied on fine-tuning diffusion models to address this issue. However, these methods typically require considerable time and computation resources, which limits their practical applicability. Recent approaches often utilize DDIM (Song, Meng, and Ermon 2020) inversion to obtain latent maps and then control the attention mechanism in diffusion models for real image editing (Mokady et al. 2023) (Cao et al. 2023). However, the consistency of the edited images depends heavily on the invertibility of the DDIM inversion process. Although some efforts have focused on optimizing this inversion (Ju et al. 2024) (Dong et al. 2023) for better results, the editing framework still relies on too many timesteps (e.g., 50 steps).
@@ 2.1 Text-to-Image Generation
Since Dosovitskiy et al. (Dosovitskiy et al. 2020) introduced the Visual Transformer (ViT) and highlighted the potential of transformer architectures for image tasks, numerous transformer-based visual applications have been developed, including high-resolution image synthesis (Esser, Rombach, and Ommer 2021). Prior to the advent of diffusion models, researchers predominantly relied on a generative adversarial network (GAN) for image synthesis (Goodfellow et al. 2014). Zhang et al.  (Zhang, Xie, and Yang 2018) developed a single-stream generator capable of producing high-resolution images, while Liang et al.  (Liang, Pei, and Lu 2020) enhanced the performance of text-to-image synthesis by incorporating Memory-Attended Text Encoder and Object-Aware Image Encoder. Subsequently, the Denoising Diffusion Probabilistic Models (DDPMs) introduced by Ho et al. (Ho, Jain, and Abbeel 2020) marked significant leaps forward, achieving breakthroughs in image quality, controllability, and diversity. The designs and applications of diffusion-based methods can be classified by tasks such as controllable generation, stylization, and quality improvement. Dhariwal et al. (Dhariwal and Nichol 2021) introduces a classifier guidance method to improve the generation quality of diffusion models, while Yang et al. (Yang et al. 2024) uses the CLIP latents to produce realistic images closely aligned with human expectations in text-to-image generation tasks. In recent text-to-image generation tasks, ControlNet (Zhang, Rao, and Agrawala 2023) allows for the integration of user-specified conditional information into the image generation process. Meanwhile, ScaleCrafter (He et al. 2023) addresses the issue of limited perception in convolutional layers during the diffusion model generation process, enabling the production of higher resolution and higher quality images. Moreover, these T2I models have also driven the development of a series of video generation and editing applications (Ma et al. 2022a, 2023, b, 2024c; Chen et al. 2024).
@@ 2.2 Interactive Image Editing
Image editing encompasses scenarios like iterative generation, collaborative creation, and image inpainting. Research has focused on decoupling high-level concepts and low-level styles within deep latent structures to improve diffusion-based models’ performance in tasks such as content editing (detail control) (Kawar et al. 2023b), style transfer (Brack et al. 2022), and textual inversion (Gal et al. 2022). Compared to other generative models, diffusion models offer enhanced controllability during the image generation process, allowing for precise manipulation of image attributes (Choi et al. 2021) (Zhang, Rao, and Agrawala 2023). These advantages enable diffusion models to achieve outstanding performance in image editing tasks. Hertz et al. (Hertz et al. 2022) introduced a framework for image editing through textual prompts, which transforms the original image into the target image by modifying, adding, and adjusting the weights of the cross-attention map. Methods like InstructPix2Pix (Brooks, Holynski, and Efros 2023) and Custom Diffusion (Kumari et al. 2023) employ user-guided approaches to achieve image editing. These techniques allow for modifications by inputting various types of guiding prompts, allowing diffusion models to swiftly adapt to new concepts. Parmar et al. (Parmar et al. 2023) employed pix2pix-zero to address the challenge of preserving the original structure while incorporating user-specified changes during image editing. Although there have been significant advancements in image editing using diffusion models, these existing attempts at image editing are still bounded by the pretrained generative power of a UNet. Compared with UNet-based diffusion models, DiT is more scalable and has more succinct architectures, while DiT’s application in image editing is still under-explored.
@@ 3 Methodology
Our proposed framework aims to achieve high-quality image editing for various sizes based on a diffusion transformer. Our method is the first editing strategy based on a pre-trained text-to-image transformer-based diffusion model, e.g., PIXART-$\alpha$  (Chen et al. 2023). With our approach, users can achieve better editing results compared to existing UNet-based methods by providing a target prompt. In this section, we first introduce the latent diffusion models and DPM inversion. Then we illustrate the superiority of transformer-based denoising in image editing tasks. Finally, we discuss the implementation details of our editing framework.
@@ 3.1 Preliminaries: Latent Diffusion Models
The Latent Diffusion Model (LDM) (Rombach et al. 2022) proposes an image generation method with a denoising process within a latent space $\mathcal{Z}$. In particular, it uses a pre-trained image encoder $\mathcal{E}$ to encode the input image $x$ into low-resolution latents $z=\mathcal{E}(x)$. During training, the model optimizes a denoising UNet $\epsilon_{\theta}$ by removing artificial noise, conditioned on both text prompt embedding $y$ and current image sample $z_{t}$, which is a noisy sample of $z_{0}$ at step $t\in[0,T]$:
@@ 3.3 Diffusion Transformer-based Image Editing
DPM-Solver inversion. As we discussed earlier, using a high-order DPM-Solver (e.g., DPM-Solver++), can effectively improve the sampling speed. To approximate the integral term $\int_{\lambda_{s}}^{\lambda_{t}}e^{-\lambda}\boldsymbol{x}_{\theta}\mathrm{d}\lambda$  in equation Eq. 4, given the $x_{t_{i-1}}$ at time $t_{t_{i-1}}$, using the Taylor expansion at $\lambda_{t_{i-1}}$, and the DPM-Solver++ can obtain a exact solution value at time $t_{i}$:
@@
The unified control of attention mechanism. In the previous work Prompt to Prompt (P2P) (Hertz et al. 2022), researchers demonstrate that the cross attention layers contain rich semantic information from prompt texts. This finding can edit images through replacing the cross attention maps between the source image and target image during the diffusion process. Specifically, the two commonly used text-guided cross attention strategies are cross attention replacement and cross-attention refinement. These two methods ensure the seamless flow of information from the target prompt to the source prompt, thereby guiding the latent map towards the desired direction.
Different from the cross attention, the self-attention mechanism in the diffusion transformer is utilized to guide the formation of image layout, a feature that cannot be accomplished by the cross-attention mechanism. As shown in Figure 3, the object and layout information from the prompt are not fully captured in the query vectors of the transformer’s shallow layers but are well-represented in the deeper layers. Moreover, with an increasing number of transformer layers, the query vectors’ ability to capture object details becomes clearer and more specific. This suggests that the transformer’s global attention mechanism is more effective at capturing long-range object information, making DiT particularly advantageous for large-scale deformation and editing of extensive images. This observation suggests that non-rigid editing of images can be achieved by controlling self attention mechanism. In MasaCtrl (Cao et al. 2023), researchers introduced the mutual self attention control mechanism. To be specific, in the early steps of diffusion, the feature in the editing steps $Q_{tar}$, $K_{tar}$, and $V_{tar}$ will be used in self attention calculation to generate an image layout closer to the target prompt, while in the later stages, the feature in the reconstruction steps ––$K_{src}$ and $V_{src}$ will be used to guide the generation of the target image layout closer to the original image.
@@
Table 1: Quantitative comparison results. We compare our model with six prior works, all implemented using official open source code.
@@ 4.1 Implementation Details
For editing tasks involving images with a scale of $512\times 512$ and larger sizes up to $1024\times 2048$, we use pre-trained models PixArt-$\alpha-\text{XL}-512\times 512$ version for the smaller scale and PixArt-$\alpha-\text{XL}-1024\times 1024-\text{MS}$ version for the larger scale (Chen et al. 2023). We conduct editing on both real and generated images. For the real image input, we use DPM-Solver inversion to get the latent noise map. We configured the DPM-Solver with 30 steps, the classifier-free guidance of 4.5, and a patches merging ratio of 0.8. All experiments were carried out using an NVIDIA Tesla A100 GPU.
@@ 4.2 Qualitative Comparison
We evaluate the qualitative performance differences between our proposed DiT4Edit editing framework and six prior baselines, all implemented using official open source code.
As shown in Figure 5, we compare our method on $512\times 512$ and $1024\times 1024$ images. The first row of the Figure 5 demonstrates our framework has the ability to generate edited images that remain consistent with the original content when editing real $512\times 512$ images, whereas existing methods often alter the background or target details of the original image. Furthermore, the second and third rows of Figure 5 illustrate our experiments with large-scale images and arbitrarily sized images—tasks that previous UNet-based methods struggled to address. The results indicate that our proposed framework effectively handles style and object shape modifications in larger images. In contrast, some state-of-the-art UNet-based methods, despite being capable of performing editing tasks, frequently result in significant alterations and damage to the background and object locations in the source image. Additionally, due to the limitations of the UNet structure, these methods typically generate target images only at a size of $512\times 512$. These findings emphasize the substantial potential of transformer-based diffusion models in large-scale image editing. We also perform the user study for comprehensive comparisons. The details of the user study can be found in the supplementary material.
@@ 4.4 Ablation Study
We perform a series of ablation studies to demonstrate the effectiveness of DPM-Solver inversion and patches merging. The results of our ablation experiments on patches merging are presented in Figure 4 and Table 2. Implementing patches merging led to a notable reduction in the editing time for large-sized images while maintaining editing quality comparable to that achieved without patches merging. This indicates that patches merging can significantly enhance the overall performance of image editing frameworks. Furthermore, the ablation experiment results for DPM-Solver and DDIM are illustrated in Figure 7. When comparing the two methods with the same number of inference steps ($T=30$), DPM Solver consistently outperformed DDIM in terms of image editing quality. This demonstrates that our use of the DPM-Solver inversion strategy allows for the generation of superior latent maps, resulting in better editing outcomes within fewer steps.